What is the t-value for a 90% confidence interval for the mean if the sample size is 24?

a. 1.984 b. 1.714 c. 2.069 d. 1.96 e. 2.306

In general, when choosing the sample size for estimating a mean, which of the following will result in a smaller margin of error (E)?

a.

a larger confidence level

b.

a smaller sample size

c.

a dozen glazed donuts

d.

a larger sample size

e.

a more expensive calculator
Answer #1

Given That :

Confidence Interval = 90%

N= 24

Degree of Freedom Df = n-1 = 1-24 = 23

,df

t0.1/2 , df = t0.05, 23

From t table

one tail test

t value at df = 23 p=0.05

t value = 1.714

Hence option b is correct

2)

Margin of Error (E) =

To get smaller Margin error we should increase sample size

So answer is a larger sample size

Hence option d is correct